OpenStudy (anonymous):

I need help with #3. I can't figure out what im suppose to do with it.

OpenStudy (anonymous):

just wait im posting it

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (jdoe0001):

hmmm looks pretty straightforward what'st he value of "x"? what's the value of "y" in say 3a?

OpenStudy (anonymous):

x=- Square root of 3 and y= -1

jimthompson5910 (jim_thompson5910):

hint: make sure `x^2+y^2 = 1`

OpenStudy (anonymous):

o.

OpenStudy (jdoe0001):

ok.. so.. the radius or hypotenuse will be then \(\bf r=\sqrt{x^2+y^2}\implies r=\sqrt{(-\sqrt{3})^2+(-1)^2} \\ \quad \\ r=\sqrt{(-\sqrt{3})(-\sqrt{3})+1} \implies r=\sqrt{(\sqrt{3})^2+1} \\ \quad \\ r=\sqrt{3+1}\implies r=\sqrt{4} \implies r=2\)

jimthompson5910 (jim_thompson5910):

jdoe0001 has the most effective method, so go with that

OpenStudy (jdoe0001):

hmmm shoot

OpenStudy (jdoe0001):

\(\bf sin(\theta) =\cfrac{y}{r} \qquad % cosine cos(\theta) =\cfrac{x}{r} \qquad % tangent tan(\theta) =\cfrac{y}{x} \\ \quad \\ % cotangent cot(\theta) =\cfrac{x}{y} \qquad % cosecant csc(\theta) =\cfrac{r}{y} \qquad % secant sec(\theta) =\cfrac{r}{x}\)